OpenStudy (anonymous):

Harold left town driving east at 55 mph. Carol left 2 hours later driving the same route at 65 mph. How many hours after Harold left will it take Carol to catch up to Harold? A. 10 B. 11 C. 12 D. 13

OpenStudy (anonymous):

let us suppose after x hours they will cross each other and try to solve

OpenStudy (anonymous):

after 2 hours he is 110 miles ahead. she travels 10mph faster so to catch him she needs to travel how many hours ?

OpenStudy (anonymous):

try this... Harold has moved for 2 hours so 110 mph. Now in the moment Carol starts. use like: Harold is 110 mph away, and he's not moving, and the Carol is moving 10 mph speed. When will they meet each other ?

OpenStudy (anonymous):

55x = 65(x-2) 55x = 65x - 130 65x - 55x = 130 10x = 130 x = 13

OpenStudy (anonymous):

sheg I think you should have: 55x(x-2)=65x IMAO. :( ... maybe I'm wrong.

OpenStudy (anonymous):

sorry 55(x-2)

OpenStudy (anonymous):

55*13 = ??? 65*11 = ???

OpenStudy (anonymous):

How many hours after Harold left will it take Carol to catch up to Harold? also remember this statement

OpenStudy (anonymous):

55(x-2) = 65x 55x-110 = 65x 55x -65x =110 -10x = 110 x = -110/10 x=-11 whoaa.. LOL

OpenStudy (anonymous):

so can it be negative

OpenStudy (anonymous):

NO.

OpenStudy (anonymous):

So you're right I guess. ;). Well done